The specification for the weight of a box of cereal is 20.8 oz ± .7 oz. The actual mean and standard deviation from a sample of 200 boxes is 20.28 oz and 0.22 oz, respectively.


What is the Cp? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

Lower specification limit=20.8-0.7=20.1

Upper specification limit=20.8+0.7=21.5

process mean=20.28

standard deviation =0.22

\\ Process\ capability =C_p\\\\ =\frac{USL-LSL}{6\sigma}\\\\ =\frac{21.5-20.1}{6(0.22)}\\\\ =\frac{1.4}{1.32}\\\\ =1.06

Lower specification limit=20.8-0.7=20.1

Upper specification limit=20.8+0.7=21.5

process mean=20.28

standard deviation =0.22

\\ Process\ capability =C_p\\\\ =\frac{USL-LSL}{6\sigma}\\\\ =\frac{21.5-20.1}{6(0.22)}\\\\ =\frac{1.4}{1.32}\\\\ =1.06